OpenStudy (1645323):

what is the conic form of this ?: y = 4x^2 - 8x + 1

OpenStudy (lucaz):

this is the standard form of a quadratic, Ax^2+Bx+C, the graph(or conic form) is a parabola

OpenStudy (1645323):

what? to get the conic form the eqaution is (y = 1/(4p)(x - h)2 + k) what is p?

OpenStudy (1645323):

@lucaz

OpenStudy (lucaz):

oh, okay, you wanted to convert the equation

OpenStudy (1645323):

yes

OpenStudy (1645323):

@lucaz ?

OpenStudy (lucaz):

first you need the vertex form y=4(x-1)^2-3 then convert to y+3=4(x-1)^2 (y+3)/4=(x-1)^2